# Optimizing importance weighting in the presence of sub-population shifts

### v1

@@ B.1 The bias-variance tradeoff
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{y_{\mathrm{te}},\bm{x}_{\mathrm{te}}}[(y_{\mathrm{te}}-f_{\hat{\bm{\theta}}_{n}(\bm{w})}(\bm{x}))^{2}|D_{\mathrm{tr}}]\right]$ $=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{y_{\mathrm{te}},\bm{x}_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|D_{\mathrm{tr}}]\right].$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}]\right]$
$=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}[y_{\mathrm{te}}^{2}-2y_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})+(\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}]\right]$
$=\mathbb{E}_{y_{\mathrm{te}}}[y_{\mathrm{te}}^{2}|g]-2\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[y_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$+\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{p(\bm{x}_{\mathrm{te}})}\left[(\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}\right]\right].$
@@
$\mathbb{E}_{y_{\mathrm{te}}}[y_{\mathrm{te}}^{2}|g]=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}+\epsilon_{\mathrm{te}})^{2}|g]=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g})^{2}|g]+\sigma^{2},$
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[y\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}+\epsilon_{\mathrm{te}})\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}|g]\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right],$
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[(\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}\right]\right]$
$=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}\left(\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right)+\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]^{2}|D_{\mathrm{tr}}\right].$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}]\right]$
$=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g})^{2}|g]-2\mathbb{E}_{\bm{x}_{\mathrm{te}}}[\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}|g]\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$+\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]^{2}|D_{\mathrm{tr}}\right]$
$+\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}\left(\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right)\right|D_{\mathrm{tr}}]+\sigma^{2}.$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}},g]\right]$
$=(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}-\bm{x}_{\mathrm{te}}^{T}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\hat{\bm{\beta}}(\bm{w})\right])^{2}+\bm{x}_{\mathrm{te}}^{T}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}(\hat{\bm{\beta}}(\bm{w})|D_{\mathrm{tr}})\right]\bm{x}_{\mathrm{te}}+\sigma^{2}_{g}.$
@@
By the law of total probability, the expected loss conditional on only $\bm{x}_{\mathrm{te}}$ can be written as a sum of of the expected losses conditional on $\bm{x}_{\mathrm{te}},g$.
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{y_{\mathrm{te}},g_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}}]\right]$
$=\sum^{G}_{g=1}p(g_{\mathrm{te}}=g)\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}},g]\right]$
$=p_{\mathrm{te}}((\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{1}-\bm{x}_{\mathrm{te}}^{T}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}[\hat{\bm{\beta}}(\bm{w})])^{2}+(1-p_{\mathrm{te}})((\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{0}-\bm{x}_{\mathrm{te}}^{T}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}[\hat{\bm{\beta}}(\bm{w})])^{2}$
$+\bm{x}_{\mathrm{te}}^{T}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}(\hat{\bm{\beta}}(\bm{w})|D_{\mathrm{tr}})\right]\bm{x}_{\mathrm{te}}+\sigma^{2}.$ (22)
@@ B.2 The asymptotic bias and variance of the coefficients
$\frac{n_{1}}{n}$ $=\frac{np_{\mathrm{tr}}}{n}=p_{\mathrm{tr}},\quad,\frac{n_{0}}{n}=\frac{n(1-p_{\mathrm{tr}})}{n}=(1-p_{\mathrm{tr}}),$ (23)
@@
$\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}=w_{1}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+w_{0}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0},\quad\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{y}_{\mathrm{tr}}=w_{1}\bm{X}_{\mathrm{tr},1}^{T}\bm{y}_{1}+w_{0}\bm{X}_{\mathrm{tr},0}^{T}\bm{y}_{0},$ (25)
and similarly $\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}=w_{1}^{2}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+w_{0}^{2}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}$, where $\bm{X}_{\mathrm{tr},1}\in\mathbb{R}^{n_{1}\times d},\bm{X}_{\mathrm{tr},0}\in\mathbb{R}^{n_{0}\times d}$ are matrices where for the observations holds that $g_{i}=1,g_{i}=0$ respectively, and similarly $\bm{y}_{1}\in\mathbb{R}^{n_{1}},\bm{y}_{0}\in\mathbb{R}^{n_{0}}$.
@@
$\left(\frac{1}{n}(w_{1}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+w_{0}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0})\right)^{-1}$ $=\left(\frac{1}{n}(w_{1}\frac{n_{1}}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+w_{0}\frac{n_{0}}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0})\right)^{-1}$
$=\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\right)^{-1}.$ (26)
And similarly,
$\frac{1}{n}\left(w_{1}^{2}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+w_{0}^{2}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\right)$ $=\frac{1}{n}\left(w_{1}^{2}\frac{n_{1}}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+w_{0}^{2}\frac{n_{0}}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\right)$
$=\left(\frac{p^{2}}{p_{\mathrm{tr}}}\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})}\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\right).$ (27)
@@
$\frac{1}{n}(w_{1}\bm{X}_{\mathrm{tr},1}^{T}\bm{y}_{1}+w_{0}\bm{X}_{\mathrm{tr},0}^{T}\bm{y}_{0})$ $=\frac{1}{n}(w_{1}\frac{n_{1}}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{y}_{1}+w_{0}\frac{n_{0}}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{y}_{0})$
$=p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{y}_{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{y}_{0}.$ (28)
Using Equation 26 and Equation 28, the WLS estimator can be re-written as
$(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{y}_{\mathrm{tr}}$ $=(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{y}_{\mathrm{tr}}$
$=\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{y}_{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{y}_{0}\right).$

### v2

@@ B.1 The bias-variance tradeoff
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{y_{\mathrm{te}},\bm{x}_{\mathrm{te}}}[(y_{\mathrm{te}}-f_{\hat{\bm{\theta}}_{n}(\bm{w})}(\bm{x}))^{2}|D_{\mathrm{tr}}]\right]$ $=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{y_{\mathrm{te}},\bm{x}_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|D_{\mathrm{tr}}]\right].$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}]\right]$
$=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}[y_{\mathrm{te}}^{2}-2y_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})+(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}]\right]$
$=\mathbb{E}_{y_{\mathrm{te}}}[y_{\mathrm{te}}^{2}|g]-2\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[y_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$+\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{p(\bm{x}_{\mathrm{te}})}\left[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}\right]\right].$
@@
$\mathbb{E}_{y_{\mathrm{te}}}[y_{\mathrm{te}}^{2}|g]=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}+\epsilon_{\mathrm{te}})^{2}|g]=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g})^{2}|g]+\sigma^{2},$
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[y\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}+\epsilon_{\mathrm{te}})\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}|g]\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right],$
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}\left[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}\right]\right]$
$=\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}\left(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right)+\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]^{2}|D_{\mathrm{tr}}\right].$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}},y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|g,D_{\mathrm{tr}}]\right]$
$=\mathbb{E}_{\bm{x}_{\mathrm{te}}}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g})^{2}|g]-2\mathbb{E}_{\bm{x}_{\mathrm{te}}}[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}|g]\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]\right]$
$+\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{\bm{x}_{\mathrm{te}}}\left[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right]^{2}|D_{\mathrm{tr}}\right]$
$+\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}\left(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w})|g,D_{\mathrm{tr}}\right)\right|D_{\mathrm{tr}}]+\sigma^{2}.$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}},g]\right]$
$=(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\hat{\bm{\beta}}(\bm{w})\right])^{2}+\bm{x}_{\mathrm{te}}^{\mathrm{T}}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}(\hat{\bm{\beta}}(\bm{w})|D_{\mathrm{tr}})\right]\bm{x}_{\mathrm{te}}+\sigma^{2}_{g}.$
@@
By the law of total probability, the expected loss conditional on only $\bm{x}_{\mathrm{te}}$ can be written as a sum of the expected losses conditional on $\bm{x}_{\mathrm{te}},g$.
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{y_{\mathrm{te}},g_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}}]\right]$
$=\sum^{G}_{g=1}p(g_{\mathrm{te}}=g)\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{y_{\mathrm{te}}}[(y_{\mathrm{te}}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}},g]\right]$
$=p_{\mathrm{te}}((\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{1}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}[\hat{\bm{\beta}}(\bm{w})])^{2}+(1-p_{\mathrm{te}})((\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{0}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}[\hat{\bm{\beta}}(\bm{w})])^{2}$
$+\bm{x}_{\mathrm{te}}^{\mathrm{T}}\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}(\hat{\bm{\beta}}(\bm{w})|D_{\mathrm{tr}})\right]\bm{x}_{\mathrm{te}}+\sigma^{2}.$ (22)
@@ B.2 The asymptotic bias and variance of the coefficients
$\frac{n_{1}}{n}$ $=\frac{np_{\mathrm{tr}}}{n}=p_{\mathrm{tr}},\quad\frac{n_{0}}{n}=\frac{n(1-p_{\mathrm{tr}})}{n}=(1-p_{\mathrm{tr}}),$ (23)
@@
$\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}=w_{1}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+w_{0}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0},\quad\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{y}_{\mathrm{tr}}=w_{1}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{y}_{1}+w_{0}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{y}_{0},$ (25)
and similarly $\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}=w_{1}^{2}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+w_{0}^{2}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}$, where $\bm{X}_{\mathrm{tr},1}\in\mathbb{R}^{n_{1}\times d},\bm{X}_{\mathrm{tr},0}\in\mathbb{R}^{n_{0}\times d}$ are matrices where for the observations holds that $g_{i}=1,g_{i}=0$ respectively, and similarly $\bm{y}_{1}\in\mathbb{R}^{n_{1}},\bm{y}_{0}\in\mathbb{R}^{n_{0}}$.
@@
$\left(\frac{1}{n}(w_{1}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+w_{0}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0})\right)^{-1}$ $=\left(\frac{1}{n}(w_{1}\frac{n_{1}}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+w_{0}\frac{n_{0}}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0})\right)^{-1}$
$=\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\right)^{-1}.$ (26)
Similarly,
$\frac{1}{n}\left(w_{1}^{2}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+w_{0}^{2}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\right)$ $=\frac{1}{n}\left(w_{1}^{2}\frac{n_{1}}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+w_{0}^{2}\frac{n_{0}}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\right)$
$=\left(\frac{p^{2}}{p_{\mathrm{tr}}}\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})}\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\right).$ (27)
@@
$\frac{1}{n}(w_{1}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{y}_{1}+w_{0}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{y}_{0})$ $=\frac{1}{n}(w_{1}\frac{n_{1}}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{y}_{1}+w_{0}\frac{n_{0}}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{y}_{0})$
$=p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{y}_{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{y}_{0}.$ (28)
Using Equations 26 and 28, the WLS estimator can be re-written as
$(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{y}_{\mathrm{tr}}$ $=(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{y}_{\mathrm{tr}}$
$=\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{y}_{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{y}_{0}\right).$